OpenStudy (anonymous):

PleeaasseeHellp! Simpler Form Of The Product: (3k+4)(2k^2-5k2)

OpenStudy (anonymous):

\[(3k+4)(2k^2-5k^2)\]?

OpenStudy (gw2011):

(3k^4+4)(2k^2-5k^2) 6k^3-15k^3+8k^2-20k^2 -9k^3-12k^2 = -3k^2(3k+4) Therefore, the answer is: -3k^2(3k+4)

OpenStudy (anonymous):

if this is actually the question, then \[(3k+4)(2k^2-5k^2)=(3k^2+4)\times -3k^2=-9k^3-12k^2\]